OpenStudy (anonymous):

Solve 2x + y = 7 x – y = –4

OpenStudy (anonymous):

y=5 x=1

OpenStudy (ash2326):

We have \[2x+y=7\] and \[x-y=-4\] We'll use substitution method here, Find x in terms of y from second equation, we have \[x-y=-4\] Add y to both the sides \[x=y-4\] Let's substitute this in first equation \[2(y-4)+y=7\] \[2y-8+y=7\] add 8 to both the sides \[3y=15\] divide both sides by 3, we get \[y=5\] now sunstitute this in the value of x we found earlier \[x=y-4\] \[x=5-4\] we get \[x=1\] So x=1 and y=5